The recent Mega Millions and Powerball lottery jackpots of more than $300 million
have raised the question: How can you make money on the lottery? We show you how,
 and warn you of things to avoid. This is not pseudo-science; all our claims are
justified with mathematics. The amazing thing is that the justifications only need
 freshman calculus, a little probability, and some economics from the 1950s, and
                    the answers can be understood by anyone.